OpenStudy (anonymous):

Can someone help me solve this Definite Integral: Integrating variable n from 0 -> (RetirementAge-CurrentAge-1). The formula to integrate is ((Salary*Deferral%)*(1+Inflation)^n)*(1+MarketRate)^(RetirementAge-CurrentAge-n). For those who are curious, this is for a Retirement Finance calculation and my hope is to convert this integral into something I can put into Excel!

OpenStudy (turingtest):

do any of the values that appear as words in this integral vary with n?

OpenStudy (turingtest):

salary, deferral, etc?

OpenStudy (anonymous):

Well, yes, in that n is used in the exponent of each part of the equation. Is that what you meant?

OpenStudy (turingtest):

no, I mean that we can rteat salary,Deferral%, market rate as constant? i.e. they are not functions of n ?

OpenStudy (turingtest):

treat ?

OpenStudy (anonymous):

That is mostly correct. Constants are Deferral%, Inflation, MarketRate, RetiremetnAge, and CurrentAge. Salary will be increasing though by the rate of Inflation each year. The first part of the equation, (Salary*Deferral%)*(1+Inflation)^n), represents the increase of Salary each year after the first by Inflation.

OpenStudy (turingtest):

well mathematically we have to either know exactly how those varaibales are functions of n, or else treat them as constant.... I don't know enough about economics to say how dependent those variables are on n and whether or not we can ignore that effect

OpenStudy (anonymous):

Treat them as constants. Only n is being integrated. Nothing else is dependent on n.

OpenStudy (anonymous):

I see. Salary is also constant. Thus, yes, aside from n, all the rest are Constants.

OpenStudy (anonymous):

My Calculus speak is not what it used to be...

OpenStudy (turingtest):

if we just say all those things are constants RetirementAge=A Current age=a Salary=S Deferral%=D Inflation=I MarketRate=r the integral is\[\int_0^{A-a-1}SD(1+I)^n(1+r)^{A-a-n}dn\]I guess...

OpenStudy (anonymous):

Yes, save that I don't know what the dn is.

OpenStudy (turingtest):

it just means that we are integrating with respect to n

OpenStudy (turingtest):

not that I know how to do this integral offhand...

OpenStudy (anonymous):

Yes, that is the correct formula. It's what I tried to type up front, but since this was my first post, I didn't realize i could do the fancy formatting you just did.

OpenStudy (anonymous):

Haha, wait, you don't know how to integrate this D:< ?

OpenStudy (turingtest):

S and D are constant, so we can tqake them out of the integral\[SD\int_0^{A-a-1}(1+I)^n(1+r)^{A-a-n}dn\]now we have two differently based exponential functions, so this is not an easy integral

OpenStudy (turingtest):

we can even pull out (1+r)^(A-a)\[SD(1+r)^{A-a}\int_0^{A-a-1}(1+I)^n(1+r)^{-n}dn\]okay now I think I can do it :)

OpenStudy (anonymous):

Very nice!

OpenStudy (turingtest):

\[SD(1+r)^{A-a}\int_0^{A-a-1}(\frac{1+I}{1+r})^ndn\]it's going to be tricky to type, but this can be integrated

OpenStudy (turingtest):

the formula to use is\[\int a^xdx=\frac{a^x}{\ln x}+C\]in your case a= that whole fraction\[\large SD(1+r)^{A-a}\int_0^{A-a-1}(\frac{1+I}{1+r})^ndn=\left .\frac{(\frac{1+I}{1+r})^n}{\ln(\frac{1+I}{1+r})}\right|_0^{A-a-1}\]

OpenStudy (turingtest):

sorry, forgot the part I took out of the integral\[ SD(1+r)^{A-a}\int_0^{A-a-1}(\frac{1+I}{1+r})^ndn= SD(1+r)^{A-a}\left .\frac{(\frac{1+I}{1+r})^n}{\ln(\frac{1+I}{1+r})}\right|_0^{A-a-1}\]

OpenStudy (turingtest):

so this is the same as\[{SD(1+r)^{A-a}\over \ln(\frac{1+I}{1+r})}\left .(\frac{1+I}{1+r})^n\right|_0^{A-a-1}\]

OpenStudy (turingtest):

evaluating n this gives\[={SD(1+r)^{A-a}\over \ln(\frac{1+I}{1+r})}[(\frac{1+I}{1+r})^{A-a-1}-1]\]not sure how you want to simplify this...

OpenStudy (turingtest):

anyway, that would be your answer if we haven't had any miscommunications along the way hope that helped!

OpenStudy (anonymous):

I think I can just sink this into excel lock/stock and it will work. I'll test and let you know!

OpenStudy (turingtest):

please do!

OpenStudy (anonymous):

Quick clarification, the part is brackets is to be multiplied by the first quotient, right?

OpenStudy (turingtest):

right

OpenStudy (turingtest):

\[{SD(1+r)^{A-a}\over \ln(\frac{1+I}{1+r})}\left[(\frac{1+I}{1+r})^{A-a-1}-1\right]\]here it is more clear

OpenStudy (turingtest):

or if you prefer this is the same as\[{SD(1+r)^{A-a}\over \ln(1+I)-\ln(1+r)}\left[(\frac{1+I}{1+r})^{A-a-1}-1\right]\]there are many ways to write the answer

OpenStudy (anonymous):

You nailed it TuringTest >:D !